OpenStudy (goformit100):

Use Euclid’s algorithm to find the HCF of 4052 and 12576.

Parth (parthkohli):

Do you know the Euclidean Algorithm?

OpenStudy (goformit100):

don't know , i read it 4 years before. So, don't remember

Parth (parthkohli):

\[12576 = 4052 \cdot 3 + 400\]\[4052 = 400 \cdot 10 + 52\]Repeat.

Parth (parthkohli):

What do you get?

OpenStudy (goformit100):

I got the answer.

OpenStudy (aravindg):

what did you get ?

OpenStudy (goformit100):

I got 4.

Parth (parthkohli):

How?

OpenStudy (goformit100):

By calculation.

Parth (parthkohli):

Oh, I meant \(12576 = 4052 \cdot 3 + 420\)